You need to use your cell phone, which broadcasts a signal at 900MHz

900MHz, but you are in an alley between two large buildings that have only a 5.0m

5.0m spacing between them. What is the angular width, in degrees, of the signal after it emerges from between the buildings?


1
Expert's answer
2021-01-22T11:41:48-0500

Two radio antennas broadcasting in phase. "\nu" = 900MHz

"dsin\theta = m\lambda \to m = 0, +1, -1, +2, -2..." 1-equation

Since the resultant wave is detected at distances much greater than d = 5m we may use 1-equation to give the directions of the intensity maxima, the values of for which the path difference is zero or a whole number of wavelengths.

The wavelength is "\lambda = \large\frac{c}{\nu}" .

From 1-equation with m = 0, and the intensity maxima are given by

"sin\theta = \large\frac{m\lambda}{d} = \large\frac{mc}{d\nu} = \frac{m*3*10^8}{5*9*10^8} = \frac{m}{15}"

"m = 0 \to \theta = 0, ... , m = 15 \to \theta = 90^o"